The period of a simple pendulum is 2.05s, What is the period of another simple pendulum which makes 300vibrations in the time it takes the first pendulum to make 200 vibration


1
Expert's answer
2021-01-10T18:25:55-0500

Let's first find the time that the first pendulum takes to make 200 vibrations:


"T_1=\dfrac{t_1}{Number\ of\ vibrations},""t_1=200T_1=200\cdot2.05\ s=410\ s."

Since we know from conditions of the question that "t_1=t_2=t", we can find the period of the second pendulum from the same formula:


"T_2=\dfrac{t_2}{Number\ of\ vibrations}=\dfrac{410\ s}{300}=1.36\ s."

Answer:

"T_2=1.36\ s."